OpenStudy (anonymous):

inverse Laplace Transform F(s)= (s^2+2)/(s^2+2s+5)

OpenStudy (anonymous):

no idea how to do this, but if you can find an answer it should match up with this http://www.wolframalpha.com/input/?i=inverse+laplace+%28s^2%2B2%29%2F%28s^2%2B2x%2B5%29

OpenStudy (anonymous):

think it is partial fractions and then a formula, but really i should just shut up

OpenStudy (anonymous):

Yeah but before the use of partial expansion you have to to make it a proper fraction.

OpenStudy (anonymous):

that i can do

OpenStudy (anonymous):

Will it be (s^2+2)/(s+1)^2+4?

OpenStudy (anonymous):

divide first \[\frac{s^2+2}{s^2+2s+5}=\frac{s^2+2s+5-2s-3}{s^2+2s+5}\] \[=1+\frac{-2x-3}{s^2+2x+5}\]

OpenStudy (anonymous):

what did you divide by?

OpenStudy (anonymous):

i just arranged it so that the power up top was less than the power in the denominator it is either that, or long division

OpenStudy (anonymous):

it is just an algebra trick so you don't have to divide

OpenStudy (anonymous):

added and subtracted what i needed to match up with the denominator

OpenStudy (anonymous):

Oh okay Thanks

OpenStudy (anonymous):

inverse laplace of 1 is the dirac delta - was that in the answer? Because if it wasn't then you may have added a solution...

OpenStudy (anonymous):

Oh I have forgotten that, thanks

OpenStudy (anonymous):

Ah, yes. I just looked at the Wolfram|Alpha link. It did indeed have the delta function.

OpenStudy (anonymous):

yeah it is in there, rest looks like greek

OpenStudy (anonymous):

Well the rest falls out of all of that fun stuff from DE class that I don't remember...

OpenStudy (anonymous):

i know you use partial fractions, and then tables i think

OpenStudy (anonymous):

but this is when i really should shut up, because i have no idea, although it seems like something i ought to learn

OpenStudy (anonymous):

I'm pretty sure OP is familiar with relations such as \[\mathcal{L}^{-1}\{\frac{1}{s^n}\} = \frac{t^{n-1}}{\Gamma (n)}\]But yeah the hard part is just the setup, the rest is looking at the table!